# Contrastive Preference Optimization: Pushing the Boundaries of LLM Performance in Machine Translation

### v3

@@ 5.2 Ablation Study
Preference Data Components: Our preference data selection involves choosing preferred and dis-preferred translations from a triplet consisting of outputs from GPT-4, ALMA, and the gold reference. In the right of Figure 4, we emphasize the significance of the data generated by both ALMA and GPT-4. The results indicate a notable decline in performance when ALMA data is excluded in the enxx direction. Conversely, omitting GPT-4 data leads to a significant performance decrease in the xxen direction. This demonstrates that data generated by both systems plays a helpful role in enhancing model performance.
5.3 Does The Quality of Dis-preferred Data Matter?
@@ 6 Conclusion
Impact Statements
This paper presents work whose goal is to advance the field of Machine Translation. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Acknowledgements
We express our profound appreciation to Hieu Hoang, Marcin Junczys-Dowmunt, Huda Khayrallah, Thamme Gowda, Vikas Raunak, Matt Post, Anoop Kunchukuttan, Roman Grundkiewicz, Philipp Koehn, Hany Hassan Awadalla, Arul Menezes, and Vishal Chowdhary for their engaging and valuable discussions that greatly enriched our work. Special thanks to Tom Kocmi for his innovative suggestion to enhance numerical data visibility using a dynamic threshold determined by estimated accuracy. Our gratitude also extends to Pushpendre Rastogi and Joey Hejna for their insightful recommendations on the CPO theory. Furthermore, we acknowledge the Unbabel Team for their valuable advice on incorporating non-COMET metrics into our analysis.
@@ D.2 Influence on Performance
Results and Analysis We show the detailed results for enxx and xxen in Table 10 and 11, respectively. The inclusion of human-labeled preference data does not significantly enhance overall translation performance. For enzh, marginal improvements are observed, though they are minimal. Conversely, for ende, a slight decline in performance is noted. In summary, the addition of human-labeled data shows no substantial difference in the enxx direction, and a minor decrease in performance for xxen on average. We hypothesize that the limited impact of these human-labeled data may stem from a high proportion of tied evaluations and potential human bias in the evaluation process. For instance, there are instances where the author consider GPT-4’s translations to be superior, while human evaluators favor those produced by Google Translate.

### v4

@@ 5.3 Ablation Study
Preference Data Components Our preference data selection involves choosing preferred and dis-preferred translations from a triplet consisting of outputs from GPT-4, ALMA, and the gold reference. In the right of Figure 4, we emphasize the significance of the data generated by both ALMA and GPT-4. The results indicate a notable decline in performance when ALMA data is excluded in the enxx direction. Conversely, omitting GPT-4 data leads to a significant performance decrease in the xxen direction. This demonstrates that data generated by both systems plays a helpful role in enhancing model performance.
@@ 6 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Translation and Large Language Model. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Acknowledgements
We express our profound appreciation to annoymous reviewers for their helpful suggestions. We also thank Tianjian Li, Hieu Hoang, Marcin Junczys-Dowmunt, Huda Khayrallah, Thamme Gowda, Vikas Raunak, Matt Post, Anoop Kunchukuttan, Roman Grundkiewicz, Philipp Koehn, Hany Hassan Awadalla, Arul Menezes, and Vishal Chowdhary for their engaging and valuable discussions that greatly enriched our work. Special thanks to Tom Kocmi for his innovative suggestion to enhance numerical data visibility using a dynamic threshold determined by estimated accuracy. Our gratitude also extends to Pushpendre Rastogi and Joey Hejna for their insightful recommendations on the CPO theory. Furthermore, we acknowledge the Unbabel Team for their valuable advice on incorporating non-COMET metrics into our analysis.
@@ D.2 Influence on Performance
Results and Analysis We show the detailed results for enxx and xxen in Table 12 and 13, respectively. The inclusion of human-labeled preference data does not significantly enhance overall translation performance. For enzh, marginal improvements are observed, though they are minimal. Conversely, for ende, a slight decline in performance is noted. In summary, the addition of human-labeled data shows no substantial difference in the enxx direction, and a minor decrease in performance for xxen on average. We hypothesize that the limited impact of these human-labeled data may stem from a high proportion of tied evaluations and potential human bias in the evaluation process. For instance, there are instances where the author consider GPT-4’s translations to be superior, while human evaluators favor those produced by Google Translate.